OpenStudy (anonymous):

If i have A(-2,4,1), B(3,5,8), and C(-4,7,3) how do i find the angle ABC?

OpenStudy (anonymous):

is their a formula?

OpenStudy (aravindg):

Simply write the equation of lines AB and BC.Then find angle between them

OpenStudy (anonymous):

how would i calculate the equations?

OpenStudy (anonymous):

we find the vectors by subtracting the corresponding co-ordinates \[ \vec{AB}=[3-(-2)]\hat{i}+(5-4)\hat{j}+(8-1)\hat{k}\\ \vec{AB}=5\hat{i}+1\hat{j}+7\hat{k} \]similarly, we find \(\vec{BC}\)

OpenStudy (aravindg):

^^

OpenStudy (anonymous):

and then i would just find the angle between them?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

and how would i do that?

OpenStudy (anonymous):

first, show me your vcetor BC

OpenStudy (anonymous):

-7i + 2j - 5k

OpenStudy (anonymous):

remember the DOT PRODUCT of vectors? using that: \[ \vec{AB}\cdot\vec{BC}=|\vec{AB}||\vec{BC}|\cos\theta\\ \cos\theta=\frac{\vec{AB}\cdot\vec{BC}}{|\vec{AB}||\vec{BC}|} \]

OpenStudy (anonymous):

good. then you'd need to find their magnitudes

OpenStudy (anonymous):

okay! i see! i get it know! thank you so much!

OpenStudy (anonymous):

you are welcome